A 230 V synchronous motor (line to line), 60 Hz, 4 poles, connected in Y has a synchronous reactance of Xs = 1.8 Ω and a negligible armature resistance. The field current flowing under these conditions is IA = 3 A. This motor is capable of absorbing a current IA = 60 A for a unit power factor (PF = 1). The linear open circuit characteristics are assumed.

(1) Is the field current in the rotor winding direct or alternating current? [1 point].

(1) Find and for these conditions. [2 points]

(2) Find the torque angle. [1 point]

(3) What is the static stability power limit (Pmax) under these conditions? [2 points] (index: Pmax = (3) / Xs)

(4) How much IA field current is needed to run the motor at 0.80 PF in advance?

1) Find V and EA
Answer #1

field. Given data: 230V Synchronus motor (line to line) 310 Frequency = 60 Hz Number of Poles - 4 Synchronus Reactance (xs) =132.79 OTC, EL-S = 132.79. 20-3(3x1-8) As we have cosp=1 > PF = 17) Ia = Ia Loº So, EL-S = 132.79 -j 5.4 (-5.4 or, EL-S= √ (1